OpenStudy (aryana_maria2323):

Please help! Will FAN AND MEDAL! Suppose the integral from 2 to 6 of g of x, dx equals 12 and the integral from 5 to 6 of g of x, dx equals negative 3, find the value of the integral from 2 to 5 of 3 times g of x, dx . a. 45 b. 15 c. 5 d. 3

OpenStudy (reemii):

\( \int_2^6 g(x) dx = \int_2^5 g(x)dx+ \int_5^6 g(x)dx \) or, in a more synthetic form: \(\int_a^b = \int_a^c + \int_c^b\). Use this to compute \(\int_2^5g(x)dx\) first.

OpenStudy (aryana_maria2323):

I don't get it?

OpenStudy (reemii):

so `the integral from 2 to 6 of g of x, dx equals 12 and the integral from 5 to 6 of g of x, dx equals negative 3` transform \[\int_2^6 g(x) dx = \int_2^5 g(x)dx+ \int_5^6 g(x)dx \qquad(1)\] into \[ 12 = \int_2^5 g(x)dx + (-3) \qquad (2)\]

OpenStudy (loser66):

my bad. I am sorry.

OpenStudy (reemii):

So that \[\int_2^5 g(x)dx = 12+3 = 15.\] Now can you compute \[ \int_2^5 3g(x)dx \quad?\]

OpenStudy (aryana_maria2323):

I replace g(x) with 15 right? but do I do it like this 3(15)?

OpenStudy (reemii):

Not that quickly. you don't know the value of \(g(x)\), only the value of the integral between 2 and 5. Instead use: \[\int_2^5 \alpha g(x) dx = \alpha \int_2^5 g(x)dx\]

OpenStudy (aryana_maria2323):

okay so now I replace g(x) with 15?

OpenStudy (reemii):

No, 15 is the value of the integral: \[ \int_2^5 3g(x)dx = 3 \underbrace{\int_2^5 g(x)dx}_{=15} = ?\]

OpenStudy (aryana_maria2323):

\[\frac{ 63g }{ 2 }\] Okay I get it, this is what I got when I plugged it into the calculator

OpenStudy (reemii):

I do \(3\times 15 = 45\). We don't know \(g(x)\) and we don't need to.

OpenStudy (aryana_maria2323):

Why don't we need to know g(x)

OpenStudy (reemii):

 the situation is as in the sketch. there's a function \(g\) there and all we know is the value of the area under the curve. total area: 12 last part: -3 => first part: 15